when three cards are selected from a deck of cards, what is the probability that all the selected cards are diamonds?
Found 2 solutions by Edwin McCravy, Alan3354:
Answer by Edwin McCravy(20063)

13 diamonds, choose 3      C(13,3)       286       11
—————————————————————  =  ————————  =  ——————— = ————— ≈ 0.0129411764705882
  52 cards, choose 3       C(52,3)      22100     850

Edwin

Answer by Alan3354(69443)

when three cards are selected from a deck of cards, what is the probability that all the selected cards are diamonds?
---------------
The 1st is 13/52 = 1/4
2nd is 12/51 = 4/17
3rd is 11/50
1*4*11/(4*17*50) = 11/850